OpenStudy (anonymous):

is there a regular polygon with an interior angle sum of 9000° ?

OpenStudy (mrnood):

exterior angles add to 360 interior angles are 180-exterior so for an n sided polygon ext = 360/n sum of int = n* (180 -360/n) = n (180n-360)/n = 180n-360 so solve for n when sum = 9000 if it is an integer then there is such a polygon...